# Beyond Prompting: An Efficient Embedding Framework for Open-Domain Question Answering

### v1

@@ 6 Conclusion
We introduce EmbQA, an embedding-level framework for open-domain QA that improves efficiency over multi-turn prompt-based systems. By refining query representations with lightweight linear layers trained via unsupervised contrastive learning, our approach reorders retrieved passages to prioritize those most likely to contain correct answers. Additionally, an exploratory embedding with an entropy-based selection mechanism enhances candidate diversity and streamlines self-verification. Experiments across multiple ODQA benchmarks, retrieval methods, and state-of-the-art LLMs show that EmbQA consistently outperforms prompt-level approaches in accuracy and efficiency.
@@ Acknowledgments
This work was supported by the UK Engineering and Physical Sciences Research Council (EPSRC) through the Turing AI Fellowship (grant nos. EP/V020579/1, EP/V020579/2). A PhD studentship from the Chinese Scholarship Council funds Zhanghao Hu. The authors also acknowledge the use of the King’s Computational Research, Engineering, and Technology Environment (CREATE) at King’s College London.
@@ A.1 Answer Candidates Generation
In Listing 1, we present the prompt $p_{can}$ which is used to generate $K$ answer candidates from the given question and $N$ retrieved passages. Here, we present the case of $K=2$.
@@ Appendix B Theoretical Discussion on Embedding Space
$\epsilon=\mathcal{\textbf{max}}_{i\neq j}|v_{i}\cdot v_{j}|^{2}$
In our work, we suppose the vectors sampled from a LLM are $v$, and the vectors introduced by injecting embeddings to the LLM are $u$. Since there are multiple tokens, including the input tokens and the token will be generated from the LLM, we mark the tokens’ vectors as $v_{i}$. Then, we have an equivalent definition under LLM setup:
$\epsilon^{\prime}=\mathcal{\textbf{max}}_{i\neq j,u}\{|v_{i}\cdot v_{j}|^{2},|v_{i},u|^{2}\}$
@@
Then, for any given injected vector $u$, the orthogonality should be decided by $|v_{i},u|^{2}$. However, the $v_{i}$ contains not only the input tokens’ representation but also the prediction of the future. It is computationally expensive to pick a reasonable $u$ after the whole decoding process. It is also meaningless to do so since we have already obtained the decoding results and no need to discuss whether the injected token embedding is efficient or not.
So, the problem is, how to estimate the potential inner product with all possible token embeddings for a given vector $u$. Here, we propose to use order statistics to find an equivalent regularization.
@@
$u_{(1)}\geq u_{(2)}\geq u_{(3)}\geq...\geq u_{(k-1)}\geq u_{(k)}$
Then we duplicate the swapping operation on $v$, which means if we swap the $i$-th dimension with $j$-th dimension on $u$, then the same operation will be deployed on $v$ as well. Here, we suppose there is a general distribution of $v$ on token embeddings and we mark the duplicating operation results as:
$v^{\prime}=\{v_{(1)},v_{(2)},v_{(3)},...,v_{(k)}\}$
@@
$\Sigma_{i=1}^{k/2}v_{(i)}u_{(i)}=\Sigma_{i=1}^{k/2}\Sigma_{j=1}^{i}\Delta_{i}v_{(j)}$
@@
$|\Sigma_{i=1}^{k}v_{(i)}u_{(i)}|^{2}<|\Sigma_{i=1}^{\frac{k}{2}}v_{(i)}u_{(i)}|^{2}+|\Sigma_{i=\frac{k}{2}+1}^{k}v_{(i)}u_{(i)}|^{2}$
@@
$|\Sigma_{i=1}^{k}v_{(i)}u_{(i)}|^{2}\leq 4\sigma_{v}^{2}\Sigma_{i=1}^{s}\Delta_{i}^{2}\mathbb{E}(e^{\frac{u_{(i)}^{2}}{2}})+2s\sigma_{v}^{2}\varepsilon$
@@
Appendix C Additional Results
C.1 EmbQA Performance Comparison between Prompt Level Rerank framework and Embedding Level Rerank Framework
Retriever& Rerank Framework HotpotQA 2Wiki NQ WebQ
EM F1 EM F1 EM F1 EM F1
BM25 25.4 37.15 16.6 21.11 26.0 32.81 22.2 31.23
+Prompt Level 39.6 52.35 19.8 28.67 39.8 51.83 36.6 50.05
+Embedding Level (Ours) 42.0 55.81 27.4 36.60 42.2 54.38 38.2 52.08
DPR 20.6 21.67 10.8 13.53 25.0 34.16 23.8 34.44
+Prompt Level 24.6 29.99 11.6 18.84 39.8 52.28 37.6 48.23
+Embedding Level (Ours) 29.8 36.34 16.8 21.03 43.0 54.36 38.0 51.95
Contriever 22.6 35.42 16.6 20.67 25.8 32.83 25.2 34.17
+Prompt Level 30.2 47.26 17.6 25.92 39.8 52.21 34.6 48.67
+Embedding Level (Ours) 36.6 52.68 26.4 34.22 42.2 53.58 36.0 49.60
Table 6: EmbQA Overall Performance on HotpotQA, 2Wiki, NQ, and WebQ datasets using LLaMA3.1.
Our experiments (Table 6) demonstrate that replacing the prompt-level rerank framework with our embedding-level alternative consistently yields superior Exact Match and F1 scores across BM25, DPR, and Contriever on HotpotQA, 2Wiki, NQ, and WebQ, thereby confirming the effectiveness and efficiency of our embedding-level rerank approach.

### v2

@@ 6 Conclusion
We introduce EmbQA, an embedding-level framework for open-domain QA that improves efficiency over multi-turn prompt-based systems. By refining query representations with lightweight linear layers trained via unsupervised contrastive learning, our approach reorders retrieved passages to prioritise those most likely to contain correct answers. Additionally, an exploratory embedding with an entropy-based selection mechanism enhances candidate diversity and streamlines self-verification. Experiments across multiple ODQA benchmarks, retrieval methods, and state-of-the-art LLMs show that EmbQA consistently outperforms prompt-level approaches in accuracy and efficiency.
@@ Acknowledgments
This work was supported in part by the UK Engineering and Physical Sciences Research Council (EPSRC) through a Turing AI Fellowship (grant no. EP/V020579/1, EP/V020579/2) and a New Horizons grant (grant no. EP/X019063/1), and KCL’s Impact Acceleration Account (grant no. EP/X525571/1). A PhD studentship from the Chinese Scholarship Council funds Zhanghao Hu. The authors also acknowledge the use of the King’s Computational Research, Engineering, and Technology Environment (CREATE) at King’s College London.
@@ A.1 Answer Candidates Generation
In Listing 1, we present the prompt $p_{can}$, which is used to generate $K$ answer candidates from the given question and $N$ retrieved passages. Here, we present the case of $K=2$.
@@ Appendix B Theoretical Discussion on Embedding Space
$\epsilon=\mathcal{\textbf{max}}_{i\neq j}|v_{i}\cdot v_{j}|^{2}.$
In our work, we suppose the vectors sampled from a LLM are $v$, and the vectors introduced by injecting embeddings to the LLM are $u$. Since there are multiple tokens, including the input tokens and the token will be generated from the LLM, we mark the tokens’ vectors as $v_{i}$. Then, we have an equivalent definition under the LLM setup:
$\epsilon^{\prime}=\mathcal{\textbf{max}}_{i\neq j,u}\{|v_{i}\cdot v_{j}|^{2},|v_{i},u|^{2}\}.$
@@
Then, for any given injected vector $u$, the orthogonality should be decided by $|v_{i},u|^{2}$. However, the $v_{i}$ contains not only the input tokens’ representation but also the prediction of the future. It is computationally expensive to pick a reasonable $u$ after the whole decoding process. It is also meaningless to do so since we have already obtained the decoding results, and no need to discuss whether the injected token embedding is efficient or not.
So, the problem is, how to estimate the potential inner product with all possible token embeddings for a given vector $u$. Here, we propose to use order statistics to find an equivalent regularisation.
@@
$u_{(1)}\geq u_{(2)}\geq u_{(3)}\geq...\geq u_{(k-1)}\geq u_{(k)}.$
Then we duplicate the swapping operation on $v$, which means if we swap the $i$-th dimension with $j$-th dimension on $u$, then the same operation will be deployed on $v$ as well. Here, we suppose there is a general distribution of $v$ on token embeddings, and we mark the duplicating operation results as:
$v^{\prime}=\{v_{(1)},v_{(2)},v_{(3)},...,v_{(k)}\},$
@@
$\Sigma_{i=1}^{k/2}v_{(i)}u_{(i)}=\Sigma_{i=1}^{k/2}\Sigma_{j=1}^{i}\Delta_{i}v_{(j)}.$
@@
$|\Sigma_{i=1}^{k}v_{(i)}u_{(i)}|^{2}<|\Sigma_{i=1}^{\frac{k}{2}}v_{(i)}u_{(i)}|^{2}+|\Sigma_{i=\frac{k}{2}+1}^{k}v_{(i)}u_{(i)}|^{2},$
@@
$|\Sigma_{i=1}^{k}v_{(i)}u_{(i)}|^{2}\leq 4\sigma_{v}^{2}\Sigma_{i=1}^{s}\Delta_{i}^{2}\mathbb{E}(e^{\frac{u_{(i)}^{2}}{2}})+2s\sigma_{v}^{2}\varepsilon.$
@@
Appendix C Overconfidence Analysis of Mistral and Qwen
To provide more detail, we break down our framework and analyse where the performance differences in Mistral and Qwen come from.
Our framework consists of two components: a retriever and a reader. As shown in Table 7 below, all three models achieve similar Top-10 retrieval passages recall for ground-truth passages across the four datasets. This indicates that the limited performance gains for Mistral and Qwen mainly arise from the reader module.
Model Framework / Avg.GT@Top-10 HotpotQA 2Wiki NQ WebQ Average
LLaMA3.1-8B-Ins Retrieval Only w/o Rerank 1.47 0.99 1.98 2.87 1.83
EmbQA (Ours) 1.87 1.49 2.55 4.31 2.56
Mistral v0.2-7B-Ins Retrieval Only w/o Rerank 1.47 0.99 1.98 2.87 1.83
EmbQA (Ours) 1.97 1.26 2.90 3.92 2.51
Qwen 2.5-7B-Ins Retrieval Only 1.47 0.99 1.98 2.87 1.83
EmbQA (Ours) 2.06 1.39 2.84 4.07 2.59
Table 7: Average ground-truth passages (GT@Top-10) across datasets and models.
Model / Avg. Candidate Entropy HotpotQA 2Wiki NQ WebQ Average
LLaMA 3.1 8B Instruct 1.06 0.98 0.95 0.93 0.98
Mistral v0.2 7B Instruct 0.17 0.19 0.20 0.20 0.19
Qwen 2.5 7B Instruct 0.15 0.20 0.14 0.17 0.17
Table 8: Average candidate entropy across datasets for different LLMs. Lower entropy indicates higher confidence in selecting relevant candidates.
We then examined answer entropy across different models and datasets. On the four context-based QA datasets used in our experiments, both Mistral and Qwen tend to produce answers with very high confidence scores (i.e., predictive probabilities). This observation is supported by the results in Table 8, where we measure the entropy of candidate phrases across datasets. Mistral and Qwen consistently yield lower entropy values, indicating a strong tendency toward overconfident predictions.
We observe that this overconfidence causes Mistral and Qwen to produce a highly skewed distribution when fed with the proposed exploratory embedding. As a result, the sum of differences we compute is significantly smaller than the theoretical estimation. This suggests that the theoretical analysis-based unified threshold used in the experimental setup for all models—based on the classical 5% significance level—may not be effective for these two models. This explains why the improvements for Mistral and Qwen were not statistically significant. Nevertheless, a key advantage of this theoretically derived threshold is that it reliably provides a performance lower bound across different models, ensuring robustness even when improvements are not statistically significant.
Moreover, we explored performance adjustments by selecting appropriate hyperparameter values for threshold T. To explore this further, we conducted ablation studies on the HotpotQA dataset, serving as a development set, with varying threshold values (Table 9) and the Contriever-based retriever. Results show that different models favour different thresholds: Mistral performs best at 0.01, while Qwen performs best at 0.005. This indicates that a fixed threshold may not be optimal across models.
Model / Threshold 0.05 0.01 0.005 0.001
Mistral v0.2 7B Ins 29.8 / 42.3 31.2 / 43.6 30.6 / 42.7 30.3 / 43.5
Qwen 2.5 7B Ins 39.0 / 50.2 39.8 / 51.5 42.0 / 52.8 40.8 / 51.9
Table 9: Effect of varying the selection threshold on EM/F1 performance for Mistral and Qwen on HotpotQA dataset.
Appendix D Further Fusion Study on Contrastive Learning and the Embedding Refinement
To clarify whether performance gains originate primarily from contrastive learning or the embedding refinement itself, we conducted an additional ablation experiment (The “New experiment" of the table 10). Specifically, we removed both the exploratory embedding and contrastive learning modules from our reranking framework, directly summing the query and candidate representations for reranking. As shown in the table, this approach yields minimal improvement, highlighting that the significant gains indeed arise from integrating contrastive learning with embedding refinement.
LLaMA 3.1 8B Instruct EM / F1 HotpotQA 2Wiki NQ WebQ Average
From Figure 2 EmbQA w/o exploratory & reranking 33.6 / 50.2 22.2 / 31.1 39.6 / 53.7 34.4 / 48.3 32.5 / 45.8
New Experiment EmbQA w/o exploratory & contrastive learning in rerank 33.8 / 50.5 22.4 / 31.4 40.2 / 53.0 34.0 / 49.1 32.6 / 46.0
From Figure 2 EmbQA w/o exploratory embedding 35.2 / 50.9 23.0 / 31.7 41.8 / 53.5 34.8 / 49.3 33.7 / 46.4
From Figure 2 EmbQA (Ours) 36.6 / 52.7 26.4 / 34.2 42.2 / 53.6 36.0 / 49.6 35.3 / 47.5
Table 10: Ablation results (EM / F1) of EmbQA under different component removal settings on LLaMA 3.1 8B Instruct.
Appendix E Detailed Comparison of Multi-turn Prompting Costs
To clarify the computational advantages of EmbQA, we provide a step-by-step breakdown of the prompting structure and cost comparison with SuRe Kim et al. (2024), as summarised below.
SuRe (7 prompts per question):
  • Step 1: Candidate Generation (1 prompt) — Generate $K=2$ initial answers from retrieved passages.
  • Step 2: Summary Generation (2 prompts) — Generate a summary for each candidate.
  • Step 3: Validity Measurement (2 prompts) — Assess factuality of each summary individually.
  • Step 4: Ranking (2 prompts) — Conduct pairwise comparisons of summaries in both orders to avoid bias.
EmbQA (2 prompts per question):
  • Step 1: Candidate Generation (1 prompt) — Generate two initial answers from retrieved passages.
  • Step 2: Re-ranking (0 prompt) — Contrastive re-ranking is performed without LLM prompting.
  • Step 3: Exploratory Embedding (1 prompt) — Inject exploration signals and generate diverse candidates.
  • Step 4: Entropy-based Selection (0 prompt) — Final answer selected based on entropy, without prompting.
Cost Scaling with Candidate Size.
SuRe scales linearly in prompt cost with the number of candidates, requiring 3 additional prompts per candidate beyond $K=2$. In contrast, EmbQA maintains a constant prompt count regardless of candidate number due to its reranking and selection mechanisms being handled without LLM queries.
Summary.
EmbQA achieves comparable or better performance using only 2 prompts per question, while SuRe requires 7. Considering both prompt count and per-step complexity, EmbQA operates at roughly one-third the runtime and output token cost of SuRe on average.
Based on the above steps, we report a detailed statistical comparison in the HotpotQA dataset as an example. Time/query means execution time per query (lower is better), and Token/query means output token requirement per query (lower is better).
Method Step Time/query (min) Token/query
SuRe Step 1 0.22 0.49k
Step 2 0.45 1.01k
Step 3 0.47 1.06k
Step 4 0.42 0.95k
Total 1.56 3.51k
EmbQA (Ours) Step 1 0.23 0.50k
Step 2 0.08 0
Step 3 0.21 0.49k
Step 4 0.01 0
Total 0.53 0.99k
Table 11: Step-wise comparison of execution time and output token cost per query between SuRe and EmbQA on HotpotQA.
Appendix F Reranking with Closed-Source API Models
Although our full EmbQA pipeline requires access to model internals (e.g., exploratory embedding and entropy-based selection), the reranking module alone can be integrated with closed-source API models. To validate this, we conduct experiments using the GPT-4o-mini model with Contriever retrieval. We compare three configurations: retrieval-only, SuRe Kim et al. (2024), and SuRe enhanced with our embedding-level reranker. Results across four ODQA datasets (Table 12) demonstrate consistent improvements in EM and F1, highlighting the generalisation potential of our reranking design.
Framework HotpotQA 2Wiki NQ WebQ Average
Retrieval-only 43.8 / 60.6 31.0 / 41.2 43.0 / 58.1 33.4 / 50.6 37.8 / 52.6
SuRe 44.8 / 59.7 34.4 / 45.1 48.4 / 60.5 38.8 / 52.6 41.6 / 54.5
SuRe + EmbQA rerank 46.9 / 62.5 35.7 / 47.5 50.9 / 62.7 42.1 / 54.8 43.9 / 56.9
Table 12: Effectiveness of the proposed embedding-level rerank module on GPT-4o-mini across four ODQA datasets.